Please help me to check my answer and correct me
Colby Limited is a manufacturing company whose total factory overhead costs fluctuate somewhat from year to year, according to the number of machine-hours worked in its production facility. These costs at high and low levels of activity over recent years are given below Level of Activity Low High Machine-hours 50,000 75,000 Total factory overhead costs $14,250,000 $17,625,000 The factory overhead costs above consist of indirect materials, rent, and maintenance. The company has analyzed these costs at the 50,000 machine-hours level of activity as follows: Indirect materials (variable) Rent (fixed) Maintenance (mixed) 5,000,000 6,000,000 3,250,000 Total factory overhead costs 14,250,000 For planning purposes, the company wants to break down the maintenance cost into its variable and fixed cost elements.
Required: 1. Estimate how much of the factory overhead cost of $17,625,000 at the high level of activity consists of maintenance cost. (Hint. First determine how much of the $17,625,000 cost consists of indirect materials and rent. Think about the behaviour of variable and fixed costs.) (Do not round intermediate calculations.) Maintenance cost 4.125,000 2. Using the high-low method, estimate a cost formula for maintenance. (Round variable cost to 1 decimal place and fixed cost to the nearest dollar amount.) 1.500000 +35 X 3. What total overhead costs would you expect the company to incur at an operating level of 70,000 machine-hours? (Do not round intermediate calculations.) Total factory overhead costs 15,450000
Answer #1
1
Total factory overhead costs 17625000
Less: Indirect materials 7500000 =5000000/50000*75000
Less: Rent 6000000
Maintenance cost 4125000
2
Maintenance cost Machine hours
High 4125000 75000
Low 3250000 50000
Difference 875000 25000
Variable cost 35 =875000/25000
Fixed cost 1500000 =4125000-(75000*35)
Y = $1500000+ $35X
3
Indirect materials 7000000 =5000000/50000*70000
Rent 6000000
Maintenance cost 3950000 =1500000+(70000*35)
Total factory overhead costs 16950000